please help me find the product of (-4/3)cibed and (-2/3)squared . for the first one I got -4^3 and the second one I got -2^2 is this correct?
Answer by checkley77(12569)

I'm afraid your answers are not correvt.
(-4/3)^3=(-64/27)
(-2/3)^2=(4/9)
(-64/27)(4/9)
-256/243 or -2^8/3^5 or 1.0535 Answer.